### v1

# 3DMambaIPF: A State Space Model for Iterative Point Cloud Filtering via Differentiable Rendering

Abstract.
Noise is an inevitable aspect of point cloud acquisition, necessitating filtering as a fundamental task within the realm of 3D vision. Existing learning-based filtering methods have shown promising capabilities on small-scale synthetic or real-world datasets. Nonetheless, the effectiveness of these methods is constrained when dealing with a substantial quantity of point clouds. This limitation primarily stems from their limited denoising capabilities for large-scale point clouds and their inclination to generate noisy outliers after denoising. The recent introduction of State Space Models (SSMs) for long sequence modeling in Natural Language Processing (NLP) presents a promising solution for handling large-scale data. Encouraged by iterative point cloud filtering methods, we introduce 3DMambaIPF, firstly incorporating Mamba (Selective SSM) architecture to sequentially handle extensive point clouds from large scenes, capitalizing on its strengths in selective input processing and long sequence modeling capabilities. Additionally, we integrate a robust and fast differentiable rendering loss to constrain the noisy points around the surface. In contrast to previous methodologies, this differentiable rendering loss enhances the visual realism of denoised geometric structures and aligns point cloud boundaries more closely with those observed in real-world objects. Extensive evaluations on datasets comprising small-scale synthetic and real-world models (typically with up to 50K points) demonstrate that our method achieves state-of-the-art results. Moreover, we showcase the superior scalability and efficiency of our method on large-scale models with about 500K points, where the majority of the existing learning-based denoising methods are unable to handle.
Point cloud filtering, denoising, state space model, differentiable rendering
1. Introduction
Point clouds, assemblies of numerous three-dimensional points denoting spatial positions, play a crucial role in representing scenes across multimedia and computer vision (CV) domains. Nevertheless, as point clouds from real-world environments are captured by devices such as LiDAR, depth cameras, or 3D scanners, noises are inevitably introduced during the scanning process (Fei et al., 2024b). The presence of these noises undermines the required data consistency and accuracy in the application scenarios of point clouds, making filtering of point clouds a fundamental task within the realm of 3D vision (Fei et al., 2022, 2023).
Due to the importance of point cloud filtering, numerous methods have been dedicated to improving the capabilities. Traditional point cloud filtering methods rely on expert priors, but they are limited by the accuracy of these priors and are not suitable for complex and sparse point clouds. In recent years, deep learning-based point cloud filtering methods have emerged prominently. These methods commonly employ iterative modules that consist of stacked encoder-decoder pairs. Such architectures leverage the output of the previous module as the input for the subsequent module, thereby demonstrating superior performance. For instance, ScoreDenoise (Luo and Hu, 2021) utilizes gradient ascent techniques and introduces an iterative point-moving method based on estimated scores. IterativePFN (de Silva Edirimuni et al., 2023) employs iterative modules and adaptive loss functions to iteratively and progressively refine noisy points toward adaptive ground truth (GT), yielding promising results on both synthetic and real-world datasets.
However, these point cloud filtering methods are only applicable to small-scale scenes and not suitable for denoising large-scale scenes. It is found experimentally that some denoising methods exhibit inadequate filtering capabilities when applied to large-scale point clouds. Furthermore, previous works only have demonstrated satisfactory performance solely in low-noise environments, but they are limited to handling details in high-noise environments. This limitation is particularly evident in the generation of unrealistic boundaries, as well as a lack of appropriate constraints. Consequently, the main challenges are twofold: (i) denoising for large-scale point clouds or complex geometries, (ii) denoising for high-noise environments.
Recently, State Space Models (SSMs), particularly structured SSMs (S4) (Gu et al., 2021a), have shown great promise in Natural Language Processing (NLP) (Lu et al., 2024). Built upon the S4 framework, Mamba (Gu and Dao, 2023) exhibits outstanding performance across many tasks in NLP and 2D CV  (He et al., 2024; Yang et al., 2024; Zhao et al., 2024; Misra et al., 2024). These achievements of Mamba can be ascribed to the linear computational complexity and its proficient long-range context learning capabilities on long sequences. Furthermore, the utilization of time-varying parameters and a hardware-aware algorithm by Mamba represents significant advancements in sequence modeling.
Considering the aforementioned challenges in point cloud filtering and the strength of Mamba for long sequence modeling, we propose 3DMambaIPF, a novel Iterative Point Clouds Filtering model utilizing the Mamba module with point rendering techniques. Specifically, 3DMambaIPF is composed of multiple iterations of Mamba-Denoise Modules, each comprising a pair of Mamba encoders and decoders. Initially, the input point cloud is divided into patches, within which a graph structure is generated based on points and their neighbors to extract position features. In the Mamba-Encoder, the Dynamic EdgeConv module encodes positional features with MLPs, which are then inputted into a Mamba module to select and generate position-dependent features from sequential inputs. The Mamba-Decoder is responsible for upsampling features into point cloud patches. During the training process, a novel rendering loss is introduced leveraging a differential point rendering technique, and the loss is computed by the disparities of rendered images of filtered point cloud and GT. Subsequently, an adaptive GT altering throughout the iterations is introduced, gradually transitioning from an initially noise-added state to the true GT. Following the completion of all iterations, a patch-stitching method is employed to reconstruct the denoised point cloud from the patches. Experimentally, 3DMambaIPF is trained using PU-Net (Yu et al., 2018) dataset and achieves state-of-the-art results compared to off-the-shelf methods. 3DMabaIPF is also able to well generalize to high-noise point clouds from the Stanford 3D Scanning Repository (Curless and Levoy, 1996; Krishnamurthy and Levoy, 1996) for synthetic objects with complex geometries. Moreover, 3DMambaIPF outperforms baseline methods in terms of visual perceptions on real-world scene datasets. In summary, our specific contributions are as follows:
  • Mamba module is integrated into the iterative filtering module, enabling modeling and filtering of long-sequence point cloud features, thereby achieving more accurate and rapid modeling of large-scale point clouds.
  • A novel, differentiable point rendering loss is integrated into 3DMambaIPF, supplanting the distance-based loss functions employed in previous methods. This enhancement endows 3DMambaIPF with the ability to handle edge denoising and makes the denoising results more realistic.
  • 3DMambaIPF surpasses baseline methods by achieving state-of-the-art results on both synthetic and real-world scene datasets, while demonstrating excellent performance even on high-noise large-scale synthetic dataset.
2. Related Work
Figure 1. Overview of 3DMambaIPF. An Encoder-Decoder-based model built on Mamba named Mamba-Denoise Module is introduced for iterative filtering. The input noisy point cloud is partitioned into patches and fed into the iterative Mamba-Denoise Module. Upon completion of the iterations, clean point cloud patches are produced as output.
Figure 2. A Gaussian noise with a variable standard deviation is added to the GT to generate the adaptive GT. As the standard deviation gradually decreases, the adaptive GT approaches the GT with each iteration. Eventually, the GT no longer changes in the final iteration.
2.1. Point Cloud Filtering Methods
Traditional point cloud filtering methods can be categorized into several main classes. Geometry-based approaches (Meek and Walton, 2000; Alexa et al., 2001, 2003; Cazals and Pouget, 2005; Huang et al., 2013; Lipman et al., 2007) involve projecting point clouds onto a surface and utilize geometric features, such as parameters derived from local surface fitting and normal estimation, to describe the shape and structure of the point cloud. Statistical-based methods (Avron et al., 2010; Sun et al., 2015; Leal et al., 2020; Zaman et al., 2017; Schall et al., 2005; Thakran and Toshniwal, 2012) employ techniques such as sparse reconstruction, density estimation and clustering analysis. Moreover, graph-based methods (Gao et al., 2018; Hu et al., 2020, 2021) utilize graph structures to represent the relationships between points. These methods leverage graph filters and graph signal processing techniques to denoise point clouds effectively. By modeling point clouds as graphs, where each point is a node and connections between points represent geometric relationships, graph-based methods can exploit the inherent connectivity and structure of the data for noise reduction. Nonetheless, traditional methods consistently prove inadequate in effectively denoising point clouds with large-scale and complex geometries.
Deep Learning-based point cloud filtering methods are undergoing significant development to address the limitations of traditional approaches. The well-known PCN (PointCleanNet) (Rakotosaona et al., 2020) model is based on PCPNet (Guerrero et al., 2018) and focuses on removing outliers and reducing noise in unordered point clouds. However, due to PCN’s denoising being performed point-wise rather than patch-wise, its efficiency is relatively low. PD-Flow (Mao et al., 2022) learns the distribution of noisy point sets and achieves denoising through noise disentanglement. However, the number of iterations needed for PD-Flow varies depending on the level of noise to achieve optimal results. GPDNet (Pistilli et al., 2020) employs graph-convolutional layers to dynamically build neighborhood graphs and utilizes the similarity of high-dimensional feature representations to construct intricate feature hierarchies. ScoreDenoise (Luo and Hu, 2021) leverages the distribution model and score estimation from noisy point clouds. Pointfilter (Zhang et al., 2020) employs an encoder-decoder architecture, where points and their neighbors are taken as input for point-wise learning. It generates a displacement vector to accomplish denoising while preserving sharp features. IterativePFN (de Silva Edirimuni et al., 2023) leverages iterative filtering modules and adaptive GT progressively yet swiftly converges noisy points onto clean surfaces, achieving superior performance. Current point cloud filtering methods are primarily designed for small-scale scenes (about 50K points) and struggle to effectively denoise point clouds with dense noise, particularly those far from clean surfaces, especially in large-scale scenes (about 500K points). Furthermore, existing methods (Zhang et al., 2020) demonstrate limited effectiveness in filtering the edges of point clouds and capturing fine-grained characteristics. Considering these challenges, our work aims to introduce a model tailored for large-scale scenes characterized by high-density noise and robust denoising capabilities at point cloud edges.
2.2. SSM-based Methods
SSMs provide a robust mathematical framework rooted in control theory and are increasingly applied in fields such as NLP and CV to model complex temporal dynamics within dynamic systems. This novel modeling approach addresses the challenge of handling long sequences, a difficulty encountered by both Transformers and RNNs. By mitigating the quadratic computational complexity of Transformers and addressing the forgetting issue of RNNs, SSMs prove highly efficient in capturing temporal dependencies within sequential data.
One noteworthy advancement in SSMs is the introduction of the HiPPO (Gu et al., 2020), which simplifies input sequence modeling by compressing inputs into a dynamic, polynomial-based representation. However, despite this innovative approach, high computational and memory footprint limit the applicability as a general sequence modeling solution (Gu et al., 2021b), failing to fully address the bottleneck issues inherent in SSMs. To this end, the S4 (Gu et al., 2021a) presents a novel parameterization technique. By integrating a low-rank structured correction, S4 achieves stable diagonalization, significantly enhancing computational efficiency through simplified Cauchy kernel operations.
Significant advancements have been made in the Mamba framework (Gu and Dao, 2023), which integrates selective SSMs into a simplified architecture. This integration allows Mamba to achieve rapid inference, linear scalability in sequence length, and state-of-the-art performance across multiple modalities, while also enhancing the selectivity of state representation. These advancements collectively demonstrate the ongoing evolution of SSMs and their increasing significance in modeling temporal dynamics across diverse domains. Methods based on Mamba have achieved state-of-the-art results in various fields, including large language models (He et al., 2024; Yang et al., 2024), medical image analysis (Liao et al., 2024; Yue and Li, 2024; Huang et al., 2024), and 3D point cloud perception (Liang et al., 2024; Zhang et al., 2024; Liu et al., 2024). Taking Consideration of the advantages of Mamba in modeling long sequences, we introduce Mamba in our work to address point cloud modeling problems in large-scale scenes.
2.3. Differentiable Rendering Techniques
Rendering techniques encompass algorithms employed to transform three-dimensional models into two-dimensional images. The differentiable rendering methods are designed for introduction into deep learning, facilitating 3D reconstruction through error backpropagation. Current differentiable rendering methods can be categorized into four types based on their geometric representations: point-based (Insafutdinov and Dosovitskiy, 2018; Yifan et al., 2019; Müller et al., 2022; Kerbl et al., 2023; Fei et al., 2024a), triangle mesh-based (Laine et al., 2020; Liu et al., 2019; Hermosilla et al., 2018), implicit function-based(Bangaru et al., 2022; Vicini et al., 2022), and volume-based(Mildenhall et al., 2021; Ren et al., 2023). In this paper, we utilize an efficient differentiable point-based rendering method (Insafutdinov and Dosovitskiy, 2018) to render both the filtered and ground-truth point clouds during training to generate the view loss, which is then propagated backward to adjust the network parameters.
3. Preliminaries
Structured State Space Models are two-stage sequence-to-sequence models that map the input $x(t)\in\mathbb{R}^{L}$ to output $y(t)\in\mathbb{R}^{L}$. Specifically, the SSM process can be described as follows:
(1) $h^{\prime}(t)=\bf{A}$ $h(t)+{\bf{B}}x(t),$
$y(t)$ $={\bf{C}}h(t),$
where $h(t)\in\mathbb{R}^{N}$ represents the latent state, and $h^{\prime}(t)\in\mathbb{R}^{N}$ is the derivative of $h(t)$. ${\bf{A}}\in\mathbb{R}^{N\times{N}}$, ${\bf{B}}\in\mathbb{R}^{N\times{1}}$, ${\bf{C}}\in\mathbb{R}^{1\times{N}}$ are the parameters. The first stage of S4 discretizes the parameter ${\bf{A}}$, ${\bf{B}}$ into ${\bf{\bar{A}}}$, ${\bf{\bar{B}}}$ with a parameter $\Delta$:
(2) ${\bf{\bar{A}}}=exp$ $(\Delta{\bf{A}}),$
${\bf{\bar{B}}}=(\Delta{\bf{A}})^{-1}(exp$ $(\Delta{\bf{A}})-I)\cdot{\Delta{\bf{B}}}.$
After discretizing, the second stage is to calculate recursively:
(3) $h_{t}={\bf{\bar{A}}}$ $h_{t-1}+{\bf{\bar{B}}}x_{t},$
$y_{t}$ $={\bf{C}}h_{t},$
with the convolution formula:
(4) ${\bf{\bar{K}}}=({\bf{C}}\bf{\bar{B}},{\bf{C}}{\bf{\bar{A}}}{\bf{\bar{B}}}$ $,...,{\bf{C}}{\bf{\bar{A}}}^{k}{\bf{\bar{B}}},...),$
$y=x$$*{\bf{\bar{K}}}.$
Figure 3. Recurrent Point Cloud Filtering of 3DMambaIPF. The Mamba Encoder comprises Dynamic EdgeConv modules and Mamba Blocks, while the Mamba Decoder consists of Mamba-Decoder Blocks and an Activated Linear module.
Mamba, as mentioned in Section 2.2, employs recurrent scans coupled with a selection mechanism to regulate the passage of specific segments within the sequence into the latent states. Specifically, by making the SSM parameters functions of the input, Mamba addresses the weakness in dealing with discrete modalities, allowing the model to selectively propagate or forget information based on the current token. Simultaneously, Mamba has devised a hardware-aware parallel algorithm to operate in recursive mode, addressing the challenge of inefficient convolution operations. The structure diagram of a single Mamba block is illustrated in Figure 4, where the overall process can be viewed as a transition from $x_{t-1}\in\mathbb{R}^{L}$ to $x_{t}\in\mathbb{R}^{L}$:
(5) ${x_{t}}^{\prime}=\text{DWConv}(\text{MLP}$ $(\text{LN}(x_{t-1}))),$
$s_{t}=\text{MLP}(\text{LN}(\text{SSM}(\sigma({x_{t}}$ )))×${}^{\prime})))\times\sigma(\text{LN}(x_{t-1}))),$
$x_{t}=s_{t}+x$ ${}_{t-1},$
where DWConv represents depth-wise convolution, LN represents layer normalization, $\sigma$ denotes activation (SiLU here).
4. Method
4.1. Overview
In this section, building on top of IterativePFN (de Silva Edirimuni et al., 2023), we will detailly introduce 3DMambaIPF which is the first Mamba framework for iteratively filtering 3D point clouds via differentiable point rendering. The recurrent point cloud filtering structure of the 3DMambaIPF model is depicted in Figure 1. 3DMambaIPF network consists of a cascade of Mamba encoders and Mamba decoders, with each pair referred to as a Mamba-Denoise Module. After receiving an extracted patch of the point cloud as input, a single Mamba-Denoise Module constructs a directed graph from the points within the patch. The directed graph is encoded into token sequences to extract features. The denoised point token sequences are then decoded, eliminating noise, and serve as the input for the next Mamba-Denoise Module. Subsequently, we will employ a differentiable 3D point rendering loss to refine the positions of noisy points. This will be achieved by comparing the rendered images from denoised point clouds with the GT. This process iterates until the final module is reached, after which the final output is a clean point cloud patch. The recurrent point cloud filtering method will be introduced in Section 4.2, the Mamba-Denoising encoder and decoder module will be introduced in Section 4.3. Section 4.4 will introduce differentiable point rendering techniques and backpropagation, while Section 4.5 will specifically describe loss function used in 3DMambaIPF.
4.2. Recurrent Point Cloud Filtering
Given a 3-dimensional clean point cloud, denoted as $X_{c}=\{x_{c1},x_{c2},...x_{cn}\}\subseteq\mathbb{R}^{3}$, each point in $X_{c}$ can be represented by three-dimensional coordinates $x_{ci}=(\alpha_{i},\beta_{i},\gamma_{i})$. To prepare the point cloud as input for the network, the K-nearest neighbors (KNN) algorithm (Peterson, 2009) is used to partition the point cloud into patches, and a directed graph is constructed within each patch. Considering a point cloud patch $X\subseteq\mathbb{K}^{3}$ which generated by KNN and a directed graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}=\{x_{1},x_{2},...x_{K}\}$ and $\mathcal{E}$ represents the edges between $x_{i}\in\mathcal{V}$ and its $k$-nearest neighbors.
[…]
Figure 4. Detailed network architectures for Dynamic EdgeConv, Mamba Block, and Mamba Decoder Block.
4.3. Mamba-Denoising Backbone
The Mamba-Denoising Backbone is the core component of 3DMambaIPF for recurrent point cloud filtering. As illustrated in Figure 3, the Mamba-Denoising backbone adopts an Encoder-Decoder structure. The encoder consists of four Dynamic EdgeConv modules and four corresponding Mamba Blocks respectively, while the decoder comprises a Mamba Decoder Block and a linear activation layer. The detailed network structure is illustrated in Figure 4.
The Dynamic EdgeConv module facilitates the conversion of input patches into a graph structure, subsequently processed through an MLP network to extract pertinent features. Initially, vertices undergo feature extraction via an MLP, capturing intrinsic characteristics. Subsequently, an additional MLP focuses on extracting positional features from vertices connected by edges to the target vertex, facilitating feature differentiation. This process culminates in concatenating the target vertex’s features with the differences of features computed from all connected vertices, effectively deriving nuanced positional features within the patch. Following the feature extraction of the point cloud within the patch using MLPs, the resultant feature sequence is sent to the Mamba Encoder. This encoder selectively processes the input feature sequence and alters it into an output feature sequence. In the Decoder module, the Mamba Decoder Block selectively processes the input feature sequence to reduce dimensionality, gradually reconstructing it into three dimensions. Linear activation is used to determine an output movement distance of a noisy point towards a clean surface after filtering. Distance values are then combined with a residual connection to the input patch, yielding the denoised point cloud patch.
4.4. Differentiable Rendering and Backpropagation
For enhanced filtering of the edges in noisy point clouds, a differentiable rendering method is devised in 3DMambaIPF, explicitly quantifying the level of noise and enabling backpropagation. To render the point clouds, 3D coordinates of the raw point cloud’s are initially converted into the standard coordinate frame via a projective transformation aligned with the camera pose. Subsequently, each discretized point is expressed as scaled Gaussian densities, yielding the occupancy map, and making the backpropagation available. Through the introduction of a differentiable ray tracing operator, these occupancies are transformed into ray termination probabilities. Ultimately, the rendered image is generated by projecting the volume onto the plane.
As depicted in Figure 5, denoised point cloud and adaptive GT are utilized to produce $K$ pairs of images with fixed camera poses using point rendering technique introduced above. These images are then used to calculate a rendering loss, which constitutes a part of the overall loss function. Since the entire rendering process is differentiable, 3DMambaIPF encourages the denoised point cloud to achieve improved geometric appearance, particularly at the boundaries, throughout the backpropagation process.
4.5. Loss Function
Our loss function comprises two components: reconstruction loss and rendering loss. The reconstruction loss aims to minimize the distance between each point in the noisy point cloud and its closest counterpart in the adaptive GT, thereby structurally denoising the point cloud. Specifically, it seeks to bring every point in the noisy point cloud closer to its corresponding (the nearest) point in the adaptive GT.
Figure 5. Diagram for rendering loss and backpropagation.
Due to the possibility of overlapping regions within each patch during patch partitioning, we introduce a patch stitching method (Zhou et al., 2022). This method weights input points based on the proximity of each point $p_{i}$ to a reference point $p_{r}$ with a Gaussian distribution. Specifically, the weight $w_{i}$ for each $p_{i}$ is calculated as Eq. 6:
(6) $w_{i}=\frac{\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})}{\sum_{i}\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})}.$
Taking reference from (de Silva Edirimuni et al., 2023), the support radius $r_{s}$ is set to $r/3$, where $r$ is the patch radius. Then, the reconstruction loss at the $t$-th iteration can be defined as Eq. 7:
(7) ${\mathcal{L}}_{Recon}(t)=\sum_{\hat{p_{i}}\in\hat{\mathcal{P}}}w_{i}\min_{{p_{ti}}\in\mathcal{P}_{t}}{{\lVert p_{ti}-\hat{p_{i}}\rVert}^{2}_{2}},$
where $\hat{\mathcal{P}}$ denotes the predicted point cloud, $\mathcal{P}_{t}$ denotes the adaptive GT at the $t$-th iteration, and $p_{i}$ represents a point in point cloud $\mathcal{P}$.
On the other hand, the rendering loss focuses on visually aligning the noisy point cloud with the adaptive GT using differentiable rendering techniques. As calculated by Eq. 8, it primarily concentrates on optimizing the boundary regions of the point cloud:
(8) ${\mathcal{L}}_{Render}(t)=\frac{1}{K}\sum_{i=1}^{K}\lvert V_{i}(\mathcal{P}_{t})-V_{i}(\hat{\mathcal{P}})\rvert,$
where $\hat{\mathcal{P}}$ denotes the predicted point cloud, $\mathcal{P}_{t}$ denotes the adaptive GT at the $t$-th iteration, $K$ is the quantity of rendered views, and $V_{i}$ represents a rendered view in a specific camera position. Loss function of one iteration is a weighted combination of the reconstruction loss and the rendering loss, and the ultimate loss is obtained by aggregating loss contributions across iterations, which are summed up as Eq. 9:
(9) ${\mathcal{L}}=\sum_{t=1}^{T}({\mathcal{L}}_{Recon}(t)+\alpha{\mathcal{L}}_{Render}(t)),$
where $T$ is number of iteration times, and weight $\alpha$ is 0.01 in experiments.
5. Experiments
5.1. Datasets and Implementation
Table 1. Results on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
Noisy 36.90 16.03 79.39 47.72 105.02 70.03 18.69 12.82 50.48 41.36 72.49 62.03
PCN 36.86 15.99 79.26 47.59 104.86 69.87 11.03 6.46 19.78 13.70 32.03 24.86
GPDNet 23.10 7.14 42.84 18.55 58.37 30.66 10.49 6.35 32.88 25.03 50.85 41.34
DMRDenoise 47.12 21.96 50.85 25.23 52.77 26.69 12.05 7.62 14.43 9.70 16.96 11.90
PDFlow 21.26 6.74 32.46 13.24 36.27 17.02 6.51 4.16 12.70 9.21 18.74 14.26
ScoreDenoise 25.22 7.54 36.83 13.80 42.32 19.04 7.16 4.00 12.89 8.33 14.45 9.58
Pointfilter 24.61 7.30 35.34 11.55 40.99 15.05 7.58 4.32 9.07 5.07 10.99 6.29
IterativePFN 20.55 5.01 30.43 8.43 33.53 10.46 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF (Ours) 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
Figure 6. Visualization Comparisons on PU-Net Dataset of 50K points with 2.5% Gaussian deviation. The coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in blue, while those with high P2F values (noisy) are portrayed in red.
[…]
Implementation. The 3DMambaIPF network is trained on NVIDIA A100 GPUs with PyTorch 1.13.1 and CUDA 11.7. The network underwent 100 epochs of training on PU-Net dataset, employing the Adam optimizer with a learning rate of 1 × $10^{-4}$. Comparative experiments are conducted with DMRDenoise (Luo and Hu, 2020), Pointfilter (Zhang et al., 2020), PD-Flow (Mao et al., 2022), ScoreDenoise (Luo and Hu, 2021) and IterativePFN (de Silva Edirimuni et al., 2023). All experimental methods (including 3DMambaIPF) are tested on the NVIDIA GeForce RTX 3090 GPU for consistency across evaluations in this paper.
5.2. Evaluation metrics
Figure 7. Visualization Comparisons on large-scale models from the Stanford 3D Scanning Repository with 5% Gaussian deviation. Coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in green, while those with high P2F values (noisy) are portrayed in purple.
Figure 8. Visualization Comparisons on Paris-rue-Madame dataset.
Table 2. Ablation results of loss function on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
IterativePFN 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF (without view loss) 20.02 4.80 30.19 8.13 32.97 10.05 5.94 2.92 7.70 4.14 9.41 5.37
3DMambaIPF (with view loss) 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
Referring to many prior notable works, we utilize Chamfer Distance (CD) and Point-to-Mesh (P2M) Distance as the evaluation metrics for our study. Notably, CD is delineated in Eq. 10, encompassing the combined sum of two components: the average distances between each point in the predicted point cloud and its nearest point in the GT point cloud, and the average distances between each point in the GT point cloud and its nearest point in the predicted point cloud.
(10) ${\mathcal{L}}_{CD}(\hat{\mathcal{P}},\mathcal{P})=\frac{1}{|\hat{\mathcal{P}}|}\sum_{{\hat{p}\in\hat{\mathcal{P}}}}\min_{{p}\in\mathcal{P}}\|\hat{p}-{p}\|+\frac{1}{|\mathcal{P}|}\sum_{{p}\in\mathcal{P}}\min_{{\hat{p}\in\hat{\mathcal{P}}}}\|{p}-\hat{p}\|,$
[…]
(11) ${\mathcal{L}}_{P2M}(\hat{\mathcal{P}},\mathcal{M})={\mathcal{L}}_{P2F}(\hat{\mathcal{P}}$ $,\mathcal{M})+{\mathcal{L}}_{F2P}(\hat{\mathcal{P}},\mathcal{M}),$
${\mathcal{L}}_{P2F}(\hat{\mathcal{P}},\mathcal{M})=\frac{1}{|\hat{\mathcal{P}}|}$ $\sum_{{\hat{p}\in\hat{\mathcal{P}}}}\min_{{f}\in\mathcal{M}}d(\hat{p},f),$
${\mathcal{L}}_{F2P}(\hat{\mathcal{P}},\mathcal{M})=\frac{1}{|\mathcal{M}|}$ $\sum_{{f}\in\mathcal{M}}\min_{{\hat{p}\in\hat{\mathcal{P}}}}d(\hat{p},f),$
where $\hat{\mathcal{P}}$ represents the predicted point cloud, $\mathcal{M}$ represents the mesh of GT point cloud, $|\mathcal{P}|$ signifies the number of points in $\mathcal{P}$, $|\mathcal{M}|$ signifies the number of faces in $\mathcal{M}$, and $d(p,f)$ denotes square distance from point $p$ to face $f$. Please note that due to the typically unequal quantities of faces and points, point-wise calculation of P2M distance is unreasonable. In the visualization of denoising results in Section 5.3 and Section 5.4, we utilize the point-wise P2F distance from Eq. 11 to visualize the denoised point cloud results.
Table 3. Ablation results of views in point rendering on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method Rendered Views 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
IterativePFN / 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF 8 20.36 4.96 30.73 8.48 34.35 10.96 6.04 2.99 8.06 4.36 10.39 6.05
3DMambaIPF 16 20.29 4.97 30.35 8.24 33.08 10.21 6.21 3.14 7.87 4.29 9.63 5.60
3DMambaIPF 24 19.96 4.82 30.26 8.14 32.85 9.95 5.96 2.97 7.81 4.26 9.56 5.53
3DMambaIPF 32 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
Table 4. Ablation results of Mamba layers on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method Mamba Layers 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
IterativePFN / 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF 1 23.35 6.76 33.24 10.91 36.42 13.17 8.15 4.71 10.82 6.53 13.23 8.20
3DMambaIPF 3 20.24 5.02 30.59 8.61 33.44 10.60 6.12 3.09 8.05 4.43 9.58 5.54
3DMambaIPF 6 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
3DMambaIPF 9 20.55 5.11 30.77 8.59 33.82 10.70 6.15 3.08 8.01 4.38 9.89 5.74
3DMambaIPF 12 20.48 5.05 30.83 8.80 33.92 10.98 6.12 3.07 8.09 4.42 9.85 5.67
Table 5. Ablation results of patch size on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
Method Patch Size 10K points 50K points
[…]
CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
IterativePFN 1000 20.55 5.01 30.43 8.43 33.53 10.46 6.05 3.02 8.03 4.36 10.15 5.88
3DMambaIPF (Ours) 1000 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
IterativePFN 2000 20.64 5.02 30.43 8.41 33.93 10.72 6.01 2.99 7.79 4.19 10.25 5.95
3DMambaIPF (Ours) 2000 19.98 4.81 30.05 8.14 32.73 10.04 5.92 2.92 7.52 4.05 8.98 5.13
5.3. Comparisons on Small-Scale Datasets
[…]
5.4. Comparisons on Large-Scale Datasets
The visualization performance of large-scale point clouds is tested on both large-scale synthetic models from the Stanford 3D Scanning Repository and Paris-rue-Madame dataset. In the visualization results of large-scale synthetic datasets, some of the baselines struggle to filter due to the presence of dense and distant noise from the surface. While PointFilter achieves filtering results where each point is close to a clean surface, it lacks significant geometric details and accurate edge structures. It is found that the incorporation of the generative model into PD-Flow results in varied denoising outcomes across different experiments on the same point cloud. Despite potential variations in outcomes, PD-Flow generally falls short in effectively denoising large-scale point clouds, particularly in densely noisy areas where it struggles to achieve significant noise reduction. ScoreDenoise, employing a patch partitioning method, exhibits gaps in results when processing large-scale point clouds. Although IterativePFN addresses this issue with patch-stitching, its performance suffers when handling point cloud edges and densely cluttered point clouds. On the contrary, the inclusion of the Mamba module enhances the processing capabilities of 3DMambaIPF when dealing with dense point clouds. Additionally, the utilization of rendering loss results in remarkable denoising effects specifically on the edges of the point cloud. In the visualization of real-world large-scale dataset (Paris-rue-Madame), when there are no additional noises introduced, all the baseline methods are capable of achieving satisfactory filtering results. 3DMambaIPF outperforms the others in terms of preserving geometric edges.
6. Ablation Study
Ablation experiments are conducted on the influencing factors of 3DMambaIPF, focusing on four aspects: i) loss function, ii) the quantity of rendered images in point rendering, iii) the number of Mamba layers, and iv) patch size. All the ablation experiments are tested on PU-Net dataset with NVIDIA GeForce RTX 3090 GPU. Initially, Table 2 validates the enhancement of the filtering capability of 3DMambaIPF across various levels of noise with the introduction of view loss. Following the previous validation step, we proceed to examine how varying numbers of rendered views affect filtering effectiveness. Table 3 suggests that generally, a greater number of views correlates with stronger denoising ability, with 32 views yielding the best results. However, employing an excessive number of views will decrease the training speed. Subsequently, ablation experiments are carried out on the number of Mamba layers. Table 4 illustrates that the denoising capability is at its peak with 6 Mamba layers. At last, the influence of patch size on denoising capability is examined in Table 5, finding that larger patch sizes are more suitable for larger scenes.
7. Conclusion
In this study, we introduce 3DMambaIPF, an iterative point cloud filtering model that utilizes the Mamba module. 3DMambaIPF not only achieves state-of-the-art results on small-scale (PU-Net) dataset, but also demonstrates superior performance on large-scale datasets, indicating its capability to address denoising challenges in extensive scenes, as validated through ablation experiments. Although 3DMambaIPF may encounter reverse-amplification issues on denoising a smooth and clean surface, its performance remains optimal.

### v2

@@
 3DMambaIPF: A State Space Model for Iterative Point Cloud Filtering via Differentiable Rendering
-Abstract.
+Abstract
 Noise is an inevitable aspect of point cloud acquisition, necessitating filtering as a fundamental task within the realm of 3D vision.
-Existing learning-based filtering methods have shown promising capabilities on small-scale synthetic or real-world datasets.
+Existing learning-based filtering methods have shown promising capabilities on commonly used datasets.
 Nonetheless, the effectiveness of these methods is constrained when dealing with a substantial quantity of point clouds.
-This limitation primarily stems from their limited denoising capabilities for large-scale point clouds and their inclination to generate noisy outliers after denoising.
-The recent introduction of State Space Models (SSMs) for long sequence modeling in Natural Language Processing (NLP) presents a promising solution for handling large-scale data.
-Encouraged by iterative point cloud filtering methods, we introduce 3DMambaIPF, firstly incorporating Mamba (Selective SSM) architecture to sequentially handle extensive point clouds from large scenes, capitalizing on its strengths in selective input processing and long sequence modeling capabilities.
-Additionally, we integrate a robust and fast differentiable rendering loss to constrain the noisy points around the surface.
+This limitation primarily stems from their limited denoising capabilities for dense and large-scale point clouds and their inclination to generate noisy outliers after denoising.
+To deal with this challenge, we introduce 3DMambaIPF, for the first time, exploiting Selective State Space Models (SSMs) architecture to handle highly-dense and large-scale point clouds, capitalizing on its strengths in selective input processing and large context modeling capabilities.
+Additionally, we present a robust and fast differentiable rendering loss to constrain the noisy points around the surface.
 In contrast to previous methodologies, this differentiable rendering loss enhances the visual realism of denoised geometric structures and aligns point cloud boundaries more closely with those observed in real-world objects.
-Extensive evaluations on datasets comprising small-scale synthetic and real-world models (typically with up to 50K points) demonstrate that our method achieves state-of-the-art results.
-Moreover, we showcase the superior scalability and efficiency of our method on large-scale models with about 500K points, where the majority of the existing learning-based denoising methods are unable to handle.
-Point cloud filtering, denoising, state space model, differentiable rendering
-1. Introduction
+Extensive evaluations on commonly used datasets (typically with up to 50K points) demonstrate that 3DMambaIPF achieves state-of-the-art results.
+Moreover, we showcase the superior scalability and efficiency of 3DMambaIPF on highly dense and large-scale point clouds with up to 500K points compared to off-the-shelf methods.
+Codehttps://github.com/TsingyuanChou/3DMambaIPF.
+Introduction
 Point clouds, assemblies of numerous three-dimensional points denoting spatial positions, play a crucial role in representing scenes across multimedia and computer vision (CV) domains.
-Nevertheless, as point clouds from real-world environments are captured by devices such as LiDAR, depth cameras, or 3D scanners, noises are inevitably introduced during the scanning process (Fei et al., 2024b).
-The presence of these noises undermines the required data consistency and accuracy in the application scenarios of point clouds, making filtering of point clouds a fundamental task within the realm of 3D vision (Fei et al., 2022, 2023).
-Due to the importance of point cloud filtering, numerous methods have been dedicated to improving the capabilities.
+Nevertheless, as point clouds from real-world environments are captured by devices such as LiDAR, depth cameras, or 3D scanners, noises are inevitably introduced during the scanning process (Fei et al.
+2024b).
+The presence of these noises undermines the required data consistency and accuracy in the application scenarios of point clouds, making filtering of point clouds a fundamental task within the realm of 3D vision (Fei et al.
+2022, 2023).
+Additionally, due to the potential large-scale nature of these real-world scans, point clouds can exhibit substantial coordinate ranges and be highly dense, containing a significant number of points.
 Traditional point cloud filtering methods rely on expert priors, but they are limited by the accuracy of these priors and are not suitable for complex and sparse point clouds.
@@
 Such architectures leverage the output of the previous module as the input for the subsequent module, thereby demonstrating superior performance.
-For instance, ScoreDenoise (Luo and Hu, 2021) utilizes gradient ascent techniques and introduces an iterative point-moving method based on estimated scores.
-IterativePFN (de Silva Edirimuni et al., 2023) employs iterative modules and adaptive loss functions to iteratively and progressively refine noisy points toward adaptive ground truth (GT), yielding promising results on both synthetic and real-world datasets.
-However, these point cloud filtering methods are only applicable to small-scale scenes and not suitable for denoising large-scale scenes.
-It is found experimentally that some denoising methods exhibit inadequate filtering capabilities when applied to large-scale point clouds.
-Furthermore, previous works only have demonstrated satisfactory performance solely in low-noise environments, but they are limited to handling details in high-noise environments.
+For instance, IterativePFN (de Silva Edirimuni et al.
+2023) employs iterative modules and adaptive loss functions to iteratively and progressively refine noisy points toward adaptive ground truth (GT), yielding promising results on both synthetic and real-world datasets.
+However, these point cloud filtering methods are only applicable to small-scale point clouds and not suitable for filtering highly dense and large-scale point clouds.
+Furthermore, previous works only have demonstrated satisfactory performance solely in low-noise environments, while limited to handling details in high-noise environments.
 This limitation is particularly evident in the generation of unrealistic boundaries, as well as a lack of appropriate constraints.
-Consequently, the main challenges are twofold: (i) denoising for large-scale point clouds or complex geometries, (ii) denoising for high-noise environments.
-Recently, State Space Models (SSMs), particularly structured SSMs (S4) (Gu et al., 2021a), have shown great promise in Natural Language Processing (NLP) (Lu et al., 2024).
-Built upon the S4 framework, Mamba (Gu and Dao, 2023) exhibits outstanding performance across many tasks in NLP and 2D CV  (He et al., 2024; Yang et al., 2024; Zhao et al., 2024; Misra et al., 2024).
-These achievements of Mamba can be ascribed to the linear computational complexity and its proficient long-range context learning capabilities on long sequences.
-Furthermore, the utilization of time-varying parameters and a hardware-aware algorithm by Mamba represents significant advancements in sequence modeling.
-Considering the aforementioned challenges in point cloud filtering and the strength of Mamba for long sequence modeling, we propose 3DMambaIPF, a novel Iterative Point Clouds Filtering model utilizing the Mamba module with point rendering techniques.
-Specifically, 3DMambaIPF is composed of multiple iterations of Mamba-Denoise Modules, each comprising a pair of Mamba encoders and decoders.
+Consequently, the main challenges are twofold: (i) denoising for highly dense and large-scale point clouds with complex geometries, (ii) denoising for high-noise environments.
+Recently, State Space Models (SSMs), particularly structured SSMs (S4) (Gu, Goel, and Ré 2021), have shown great promise in Natural Language Processing (NLP) (Lu et al.
+2024).
+Built upon the S4 framework, Mamba (Gu and Dao 2023) exhibits outstanding performance across many tasks in NLP and 2D CV  (He et al.
+2024; Yang et al.
+2024; Zhao et al.
+2024; Misra, Gala, and Orvieto 2024).
+The remarkable achievements of Mamba are a testament to its linear computational complexity and exceptional long-range context learning capabilities on extensive sequences.
+Moreover, the integration of time-varying parameters and a hardware-aware algorithm within Mamba marks a significant leap forward in the field of sequence modeling.
+However, exploiting SSM models for point cloud filtering has yet to be fully explored.
+Direct application of Mamba may not fully address the challenges involved.
+Therefore, we propose 3DMambaIPF, a novel Iterative Point Clouds Filtering model utilizing the Mamba module with differentiable point rendering techniques.
+Specifically, 3DMambaIPF is composed of multiple iterations of Mamba-Denoising Modules, each comprising a pair of Mamba encoders and decoders.
 Initially, the input point cloud is divided into patches, within which a graph structure is generated based on points and their neighbors to extract position features.
-In the Mamba-Encoder, the Dynamic EdgeConv module encodes positional features with MLPs, which are then inputted into a Mamba module to select and generate position-dependent features from sequential inputs.
-The Mamba-Decoder is responsible for upsampling features into point cloud patches.
+In the Mamba-based Encoder, the Dynamic EdgeConv module encodes positional features with MLPs, which are then fed into a Mamba module to select and generate position-dependent features from sequential inputs.
+The Mamba-based Decoder is responsible for upsampling features into point cloud patches.
 During the training process, a novel rendering loss is introduced leveraging a differential point rendering technique, and the loss is computed by the disparities of rendered images of filtered point cloud and GT.
 Subsequently, an adaptive GT altering throughout the iterations is introduced, gradually transitioning from an initially noise-added state to the true GT.
-Following the completion of all iterations, a patch-stitching method is employed to reconstruct the denoised point cloud from the patches.
-Experimentally, 3DMambaIPF is trained using PU-Net (Yu et al., 2018) dataset and achieves state-of-the-art results compared to off-the-shelf methods.
-3DMabaIPF is also able to well generalize to high-noise point clouds from the Stanford 3D Scanning Repository (Curless and Levoy, 1996; Krishnamurthy and Levoy, 1996) for synthetic objects with complex geometries.
+Following the completion of all iterations, a patch-stitching method is employed to reconstruct the filtered point cloud from the patches.
+Experimentally, 3DMambaIPF is trained using PU-Net (Yu et al.
+2018) dataset and achieves state-of-the-art results compared to off-the-shelf methods.
+3DMabaIPF is also able to well generalize to high-noise point clouds from the Stanford 3D Scanning Repository (Curless and Levoy 1996; Krishnamurthy and Levoy 1996) for synthetic objects with complex geometries.
 Moreover, 3DMambaIPF outperforms baseline methods in terms of visual perceptions on real-world scene datasets.
-In summary, our specific contributions are as follows:
-  • Mamba module is integrated into the iterative filtering module, enabling modeling and filtering of long-sequence point cloud features, thereby achieving more accurate and rapid modeling of large-scale point clouds.
-  • A novel, differentiable point rendering loss is integrated into 3DMambaIPF, supplanting the distance-based loss functions employed in previous methods.
-This enhancement endows 3DMambaIPF with the ability to handle edge denoising and makes the denoising results more realistic.
-  • 3DMambaIPF surpasses baseline methods by achieving state-of-the-art results on both synthetic and real-world scene datasets, while demonstrating excellent performance even on high-noise large-scale synthetic dataset.
-2. Related Work
-Figure 1.
-Overview of 3DMambaIPF.
-An Encoder-Decoder-based model built on Mamba named Mamba-Denoise Module is introduced for iterative filtering.
-The input noisy point cloud is partitioned into patches and fed into the iterative Mamba-Denoise Module.
+In summary, our main contributions are as follows:
+  • A novel iterative Mamba-based backbone is designed for point cloud filtering, allowing for the effective modeling of long-sequence point cloud features.
+This approach enhances the accuracy and speed of highly dense and large-scale point cloud processing.
+  • A well-designed differentiable point rendering loss is introduced into 3DMambaIPF, supplanting the distance-based loss functions.
+This differentiable rendering loss endows 3DMambaIPF with the ability to handle edge denoising and makes the denoising results more realistic.
+  • 3DMambaIPF surpasses existing methods by achieving state-of-the-art results on the commonly used PU-Net dataset, while demonstrating excellent performance even on highly-dense and large-scale datasets with high noise added.
+Related Work
+Figure 1: Overview of 3DMambaIPF.
+An Encoder-Decoder-based model built on Mamba named Mamba-Denoising Module is introduced for iterative filtering.
+The input noisy point cloud is partitioned into patches and fed into the iterative Mamba-Denoising Module.
 Upon completion of the iterations, clean point cloud patches are produced as output.
-Figure 2.
-A Gaussian noise with a variable standard deviation is added to the GT to generate the adaptive GT.
-As the standard deviation gradually decreases, the adaptive GT approaches the GT with each iteration.
-Eventually, the GT no longer changes in the final iteration.
-2.1. Point Cloud Filtering Methods
-Traditional point cloud filtering methods can be categorized into several main classes.
-Geometry-based approaches (Meek and Walton, 2000; Alexa et al., 2001, 2003; Cazals and Pouget, 2005; Huang et al., 2013; Lipman et al., 2007) involve projecting point clouds onto a surface and utilize geometric features, such as parameters derived from local surface fitting and normal estimation, to describe the shape and structure of the point cloud.
-Statistical-based methods (Avron et al., 2010; Sun et al., 2015; Leal et al., 2020; Zaman et al., 2017; Schall et al., 2005; Thakran and Toshniwal, 2012) employ techniques such as sparse reconstruction, density estimation and clustering analysis.
-Moreover, graph-based methods (Gao et al., 2018; Hu et al., 2020, 2021) utilize graph structures to represent the relationships between points.
-These methods leverage graph filters and graph signal processing techniques to denoise point clouds effectively.
-By modeling point clouds as graphs, where each point is a node and connections between points represent geometric relationships, graph-based methods can exploit the inherent connectivity and structure of the data for noise reduction.
-Nonetheless, traditional methods consistently prove inadequate in effectively denoising point clouds with large-scale and complex geometries.
+To enhance the filtering of noisy points around the surface, a differentiable rendering method is introduced.
+The rendering loss and reconstruction loss are jointly backpropagated to update the parameters of 3DMambaIPF.
+Deep Learning-based Point Cloud Filtering Methods.
 Deep Learning-based point cloud filtering methods are undergoing significant development to address the limitations of traditional approaches.
-The well-known PCN (PointCleanNet) (Rakotosaona et al., 2020) model is based on PCPNet (Guerrero et al., 2018) and focuses on removing outliers and reducing noise in unordered point clouds.
+The well-known PCN (PointCleanNet) (Rakotosaona et al.
+2020) model is based on PCPNet (Guerrero et al.
+2018) and focuses on removing outliers and reducing noise in unordered point clouds.
 However, due to PCN’s denoising being performed point-wise rather than patch-wise, its efficiency is relatively low.
-PD-Flow (Mao et al., 2022) learns the distribution of noisy point sets and achieves denoising through noise disentanglement.
+PD-Flow (Mao et al.
+2022) learns the distribution of noisy point sets and achieves denoising through noise disentanglement.
 However, the number of iterations needed for PD-Flow varies depending on the level of noise to achieve optimal results.
-GPDNet (Pistilli et al., 2020) employs graph-convolutional layers to dynamically build neighborhood graphs and utilizes the similarity of high-dimensional feature representations to construct intricate feature hierarchies.
-ScoreDenoise (Luo and Hu, 2021) leverages the distribution model and score estimation from noisy point clouds.
-Pointfilter (Zhang et al., 2020) employs an encoder-decoder architecture, where points and their neighbors are taken as input for point-wise learning.
+GPDNet (Pistilli et al.
+2020) employs graph-convolutional layers to dynamically build neighborhood graphs and utilizes the similarity of high-dimensional feature representations to construct intricate feature hierarchies.
+ScoreDenoise (Luo and Hu 2021) leverages the distribution model and score estimation from noisy point clouds.
+Pointfilter (Zhang et al.
+2020) employs an encoder-decoder architecture, where points and their neighbors are taken as input for point-wise learning.
 It generates a displacement vector to accomplish denoising while preserving sharp features.
-IterativePFN (de Silva Edirimuni et al., 2023) leverages iterative filtering modules and adaptive GT progressively yet swiftly converges noisy points onto clean surfaces, achieving superior performance.
-Current point cloud filtering methods are primarily designed for small-scale scenes (about 50K points) and struggle to effectively denoise point clouds with dense noise, particularly those far from clean surfaces, especially in large-scale scenes (about 500K points).
-Furthermore, existing methods (Zhang et al., 2020) demonstrate limited effectiveness in filtering the edges of point clouds and capturing fine-grained characteristics.
-Considering these challenges, our work aims to introduce a model tailored for large-scale scenes characterized by high-density noise and robust denoising capabilities at point cloud edges.
-2.2. SSM-based Methods
+IterativePFN (de Silva Edirimuni et al.
+2023) leverages iterative filtering modules and adaptive GT progressively yet swiftly converges noisy points onto clean surfaces, achieving superior performance.
+Current point cloud filtering methods are primarily designed for sparse and small-scale datasets (up to 50K points) and struggle to effectively filter dense point clouds (approximately 500K points) with noise far from clean surfaces, particularly in large-scale point clouds.
+Furthermore, existing methods (Zhang et al.
+2020) demonstrate limited effectiveness in filtering the edges of point clouds and capturing fine-grained characteristics.
+Considering these challenges, our work aims to introduce a model tailored for highly dense and large-scale point clouds characterized by high-density noise and robust filtering capabilities at point cloud edges.
+SSM-based Methods.
 SSMs provide a robust mathematical framework rooted in control theory and are increasingly applied in fields such as NLP and CV to model complex temporal dynamics within dynamic systems.
@@
 By mitigating the quadratic computational complexity of Transformers and addressing the forgetting issue of RNNs, SSMs prove highly efficient in capturing temporal dependencies within sequential data.
-One noteworthy advancement in SSMs is the introduction of the HiPPO (Gu et al., 2020), which simplifies input sequence modeling by compressing inputs into a dynamic, polynomial-based representation.
-However, despite this innovative approach, high computational and memory footprint limit the applicability as a general sequence modeling solution (Gu et al., 2021b), failing to fully address the bottleneck issues inherent in SSMs.
-To this end, the S4 (Gu et al., 2021a) presents a novel parameterization technique.
-By integrating a low-rank structured correction, S4 achieves stable diagonalization, significantly enhancing computational efficiency through simplified Cauchy kernel operations.
-Significant advancements have been made in the Mamba framework (Gu and Dao, 2023), which integrates selective SSMs into a simplified architecture.
+Significant advancements have been made in the Mamba framework (Gu and Dao 2023), which integrates selective SSMs into a simplified architecture.
 This integration allows Mamba to achieve rapid inference, linear scalability in sequence length, and state-of-the-art performance across multiple modalities, while also enhancing the selectivity of state representation.
 These advancements collectively demonstrate the ongoing evolution of SSMs and their increasing significance in modeling temporal dynamics across diverse domains.
-Methods based on Mamba have achieved state-of-the-art results in various fields, including large language models (He et al., 2024; Yang et al., 2024), medical image analysis (Liao et al., 2024; Yue and Li, 2024; Huang et al., 2024), and 3D point cloud perception (Liang et al., 2024; Zhang et al., 2024; Liu et al., 2024).
-Taking Consideration of the advantages of Mamba in modeling long sequences, we introduce Mamba in our work to address point cloud modeling problems in large-scale scenes.
-2.3. Differentiable Rendering Techniques
+Methods based on Mamba have achieved state-of-the-art results in various fields, including large language models (He et al.
+2024), medical image analysis (Liao et al.
+2024), and 3D point cloud perception (Liang et al.
+2024).
+Taking Consideration of the advantages of Mamba in modeling long sequences, we introduce Mamba in our work to address highly-dense and large-scale point cloud modeling problems.
+Differentiable Rendering Techniques.
 Rendering techniques encompass algorithms employed to transform three-dimensional models into two-dimensional images.
 The differentiable rendering methods are designed for introduction into deep learning, facilitating 3D reconstruction through error backpropagation.
-Current differentiable rendering methods can be categorized into four types based on their geometric representations: point-based (Insafutdinov and Dosovitskiy, 2018; Yifan et al., 2019; Müller et al., 2022; Kerbl et al., 2023; Fei et al., 2024a), triangle mesh-based (Laine et al., 2020; Liu et al., 2019; Hermosilla et al., 2018), implicit function-based(Bangaru et al., 2022; Vicini et al., 2022), and volume-based(Mildenhall et al., 2021; Ren et al., 2023).
-In this paper, we utilize an efficient differentiable point-based rendering method (Insafutdinov and Dosovitskiy, 2018) to render both the filtered and ground-truth point clouds during training to generate the view loss, which is then propagated backward to adjust the network parameters.
-3. Preliminaries
-Structured State Space Models are two-stage sequence-to-sequence models that map the input $x(t)\in\mathbb{R}^{L}$ to output $y(t)\in\mathbb{R}^{L}$.
-Specifically, the SSM process can be described as follows:
-(1) $h^{\prime}(t)=\bf{A}$ $h(t)+{\bf{B}}x(t),$
-$y(t)$ $={\bf{C}}h(t),$
-where $h(t)\in\mathbb{R}^{N}$ represents the latent state, and $h^{\prime}(t)\in\mathbb{R}^{N}$ is the derivative of $h(t)$.
-${\bf{A}}\in\mathbb{R}^{N\times{N}}$, ${\bf{B}}\in\mathbb{R}^{N\times{1}}$, ${\bf{C}}\in\mathbb{R}^{1\times{N}}$ are the parameters.
-The first stage of S4 discretizes the parameter ${\bf{A}}$, ${\bf{B}}$ into ${\bf{\bar{A}}}$, ${\bf{\bar{B}}}$ with a parameter $\Delta$:
-(2) ${\bf{\bar{A}}}=exp$ $(\Delta{\bf{A}}),$
-${\bf{\bar{B}}}=(\Delta{\bf{A}})^{-1}(exp$ $(\Delta{\bf{A}})-I)\cdot{\Delta{\bf{B}}}.$
-After discretizing, the second stage is to calculate recursively:
-(3) $h_{t}={\bf{\bar{A}}}$ $h_{t-1}+{\bf{\bar{B}}}x_{t},$
-$y_{t}$ $={\bf{C}}h_{t},$
-with the convolution formula:
-(4) ${\bf{\bar{K}}}=({\bf{C}}\bf{\bar{B}},{\bf{C}}{\bf{\bar{A}}}{\bf{\bar{B}}}$ $,...,{\bf{C}}{\bf{\bar{A}}}^{k}{\bf{\bar{B}}},...),$
-$y=x$$*{\bf{\bar{K}}}.$
-Figure 3.
-Recurrent Point Cloud Filtering of 3DMambaIPF.
-The Mamba Encoder comprises Dynamic EdgeConv modules and Mamba Blocks, while the Mamba Decoder consists of Mamba-Decoder Blocks and an Activated Linear module.
-Mamba, as mentioned in Section 2.2, employs recurrent scans coupled with a selection mechanism to regulate the passage of specific segments within the sequence into the latent states.
-Specifically, by making the SSM parameters functions of the input, Mamba addresses the weakness in dealing with discrete modalities, allowing the model to selectively propagate or forget information based on the current token.
-Simultaneously, Mamba has devised a hardware-aware parallel algorithm to operate in recursive mode, addressing the challenge of inefficient convolution operations.
-The structure diagram of a single Mamba block is illustrated in Figure 4, where the overall process can be viewed as a transition from $x_{t-1}\in\mathbb{R}^{L}$ to $x_{t}\in\mathbb{R}^{L}$:
-(5) ${x_{t}}^{\prime}=\text{DWConv}(\text{MLP}$ $(\text{LN}(x_{t-1}))),$
-$s_{t}=\text{MLP}(\text{LN}(\text{SSM}(\sigma({x_{t}}$ )))×${}^{\prime})))\times\sigma(\text{LN}(x_{t-1}))),$
-$x_{t}=s_{t}+x$ ${}_{t-1},$
-where DWConv represents depth-wise convolution, LN represents layer normalization, $\sigma$ denotes activation (SiLU here).
-4. Method
-4.1. Overview
-In this section, building on top of IterativePFN (de Silva Edirimuni et al., 2023), we will detailly introduce 3DMambaIPF which is the first Mamba framework for iteratively filtering 3D point clouds via differentiable point rendering.
-The recurrent point cloud filtering structure of the 3DMambaIPF model is depicted in Figure 1.
-3DMambaIPF network consists of a cascade of Mamba encoders and Mamba decoders, with each pair referred to as a Mamba-Denoise Module.
-After receiving an extracted patch of the point cloud as input, a single Mamba-Denoise Module constructs a directed graph from the points within the patch.
+Current differentiable rendering methods can be categorized into four types based on their geometric representations: point-based (Yifan et al.
+2019; Müller, Weinmann, and Klein 2022; Kerbl et al.
+2023; Fei et al.
+2024a), triangle mesh-based (Laine et al.
+2020; Liu et al.
+2019), implicit function-based(Bangaru et al.
+2022; Vicini, Speierer, and Jakob 2022), and volume-based(Mildenhall et al.
+2021; Ren et al.
+2023).
+In this paper, we propose an efficient differentiable rendering method based on point-based rendering (Insafutdinov and Dosovitskiy 2018) to render both the filtered and GT point clouds during training to generate the view loss, which is then propagated backward to adjust the network parameters.
+Method
+Overview
+Figure 2: A Gaussian noise with a variable standard deviation is added to the GT to generate the adaptive GT.
+As the standard deviation gradually decreases, the adaptive GT approaches the GT with each iteration.
+Eventually, the GT no longer changes in the final iteration.
+Figure 3: Mamba-Denoising Module for iterative point cloud filtering of 3DMambaIPF.
+The Mamba-based Encoder comprises Dynamic EdgeConv modules and Mamba Blocks, while the Mamba-based Decoder consists of Mamba Decoder Blocks and an Activated Linear module.
+Figure 4: Detailed network architectures for Dynamic EdgeConv, Mamba Block, and Mamba Decoder Block.
+The overview of 3DMambaIPF is depicted in Figure 1, which utilizes a novel iterative Mamba-Denosing backbone and a differentiable point rendering technique.
+Mamba-Denoising backbone consists of a cascade of Mamba-based encoders and decoders, with each pair referred to as a Mamba-Denoising Module.
+After receiving an extracted patch of the point cloud as input, a single Mamba-Denoising Module constructs a directed graph from the points within the patch.
 The directed graph is encoded into token sequences to extract features.
-The denoised point token sequences are then decoded, eliminating noise, and serve as the input for the next Mamba-Denoise Module.
-Subsequently, we will employ a differentiable 3D point rendering loss to refine the positions of noisy points.
-This will be achieved by comparing the rendered images from denoised point clouds with the GT.
-This process iterates until the final module is reached, after which the final output is a clean point cloud patch.
-The recurrent point cloud filtering method will be introduced in Section 4.2, the Mamba-Denoising encoder and decoder module will be introduced in Section 4.3.
-Section 4.4 will introduce differentiable point rendering techniques and backpropagation, while Section 4.5 will specifically describe loss function used in 3DMambaIPF.
-4.2. Recurrent Point Cloud Filtering
-Given a 3-dimensional clean point cloud, denoted as $X_{c}=\{x_{c1},x_{c2},...x_{cn}\}\subseteq\mathbb{R}^{3}$, each point in $X_{c}$ can be represented by three-dimensional coordinates $x_{ci}=(\alpha_{i},\beta_{i},\gamma_{i})$.
-To prepare the point cloud as input for the network, the K-nearest neighbors (KNN) algorithm (Peterson, 2009) is used to partition the point cloud into patches, and a directed graph is constructed within each patch.
+The denoised point token sequences are then decoded, eliminating noise, and serve as the input for the next Mamba-Denoising Module.
+Specifically, given a 3-dimensional clean point cloud, denoted as $X_{c}=\{x_{c1},x_{c2},...x_{cn}\}\subseteq\mathbb{R}^{3}$, each point in $X_{c}$ can be represented by 3-dimensional coordinates $x_{ci}=(\alpha_{i},\beta_{i},\gamma_{i})$.
+To prepare the point cloud as input for 3DMambaIPF, the K-nearest neighbors (KNN) algorithm (Peterson 2009) is used to partition the point cloud into patches, and a directed graph is constructed within each patch.
 Considering a point cloud patch $X\subseteq\mathbb{K}^{3}$ which generated by KNN and a directed graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}=\{x_{1},x_{2},...x_{K}\}$ and $\mathcal{E}$ represents the edges between $x_{i}\in\mathcal{V}$ and its $k$-nearest neighbors.
 […]
-Figure 4. Detailed network architectures for Dynamic EdgeConv, Mamba Block, and Mamba Decoder Block.
-4.3. Mamba-Denoising Backbone
-The Mamba-Denoising Backbone is the core component of 3DMambaIPF for recurrent point cloud filtering.
+Subsequently, a differentiable 3D point rendering loss is employed to refine the positions of noisy points, especially around the surface of the point clouds, by comparing the rendered images from filtered point clouds with the adaptive GT.
+Mamba-Denoising Backbone
+The Mamba-Denoising Backbone is the core component of 3DMambaIPF for iterative point cloud filtering.
 As illustrated in Figure 3, the Mamba-Denoising backbone adopts an Encoder-Decoder structure.
@@
 Subsequently, an additional MLP focuses on extracting positional features from vertices connected by edges to the target vertex, facilitating feature differentiation.
+Specifically, vertex features are updated as ${h_{i}}^{l+1}=f({h_{i}}^{l})+\sum_{i,j}g(\text{concat}({h_{i}}^{l},{h_{j}}^{l})-{h_{i}}^{l}),$ where $h^{l}$ denotes to features at layer $l$, $f(\cdot)$, $g(\cdot)$ denote one-layer or two-layers MLPs, $i$ is a vertex on the graph, $(i,j)$ forms an edge.
 This process culminates in concatenating the target vertex’s features with the differences of features computed from all connected vertices, effectively deriving nuanced positional features within the patch.
@@
 In the Decoder module, the Mamba Decoder Block selectively processes the input feature sequence to reduce dimensionality, gradually reconstructing it into three dimensions.
-Linear activation is used to determine an output movement distance of a noisy point towards a clean surface after filtering.
-Distance values are then combined with a residual connection to the input patch, yielding the denoised point cloud patch.
-4.4. Differentiable Rendering and Backpropagation
-For enhanced filtering of the edges in noisy point clouds, a differentiable rendering method is devised in 3DMambaIPF, explicitly quantifying the level of noise and enabling backpropagation.
-To render the point clouds, 3D coordinates of the raw point cloud’s are initially converted into the standard coordinate frame via a projective transformation aligned with the camera pose.
+Linear activation is used to determine the output movement distance of a noisy point towards a clean surface after filtering.
+Utilizing a residual connection, the noisy points within the input patch are adjusted under the guidance of the movement distance, resulting in a denoised patch.
+Differentiable Rendering and Backpropagation
+For enhanced filtering of the noisy points around the surface, we propose a differentiable rendering method and integrate it in 3DMambaIPF to explicitly quantify the level of noise and enable backpropagation.
+To render the point clouds, the 3D coordinates of the raw point clouds are first converted into the standard coordinate frame via a projective transformation aligned with the camera pose.
 Subsequently, each discretized point is expressed as scaled Gaussian densities, yielding the occupancy map, and making the backpropagation available.
@@
 Ultimately, the rendered image is generated by projecting the volume onto the plane.
-As depicted in Figure 5, denoised point cloud and adaptive GT are utilized to produce $K$ pairs of images with fixed camera poses using point rendering technique introduced above.
-These images are then used to calculate a rendering loss, which constitutes a part of the overall loss function.
-Since the entire rendering process is differentiable, 3DMambaIPF encourages the denoised point cloud to achieve improved geometric appearance, particularly at the boundaries, throughout the backpropagation process.
-4.5. Loss Function
+Denoised point cloud and adaptive GT are utilized to produce $K$ pairs of images with fixed camera poses using point rendering technique introduced above.
+Our differentiable rendering method employs three projection planes, allowing for the rendering of images with greater detail and coverage of the overall shapes, rather than projecting them onto a single plane.
+The rendering loss is defined in Equation 1.
+It aims to characterize the local distribution of points and penalize noise distribution, encouraging the denoised point cloud to achieve an improved geometric appearance, particularly around the surfaces of the point cloud, throughout the backpropagation process:
+${\mathcal{L}}_{Render}(t)=\frac{1}{K}\sum_{i=1}^{K}\lvert V_{i}(\mathcal{P}_{t})-V_{i}(\hat{\mathcal{P}})\rvert,$ (1)
+where $\hat{\mathcal{P}}$ denotes the predicted point cloud, $\mathcal{P}_{t}$ denotes the adaptive GT at the $t$-th iteration, $K$ is the quantity of rendered views, and $V_{i}$ represents a rendered view in a specific camera position.
+Loss Function
 Our loss function comprises two components: reconstruction loss and rendering loss.
@@
 Specifically, it seeks to bring every point in the noisy point cloud closer to its corresponding (the nearest) point in the adaptive GT.
-Figure 5. Diagram for rendering loss and backpropagation.
-Due to the possibility of overlapping regions within each patch during patch partitioning, we introduce a patch stitching method (Zhou et al., 2022).
+Due to the possibility of overlapping regions within each patch during patch partitioning, we introduce a patch stitching method (Zhou et al.
+2022).
 This method weights input points based on the proximity of each point $p_{i}$ to a reference point $p_{r}$ with a Gaussian distribution.
-Specifically, the weight $w_{i}$ for each $p_{i}$ is calculated as Eq.
-6:
-(6) $w_{i}=\frac{\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})}{\sum_{i}\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})}.$
-Taking reference from (de Silva Edirimuni et al., 2023), the support radius $r_{s}$ is set to $r/3$, where $r$ is the patch radius.
-Then, the reconstruction loss at the $t$-th iteration can be defined as Eq.
-7:
-(7) ${\mathcal{L}}_{Recon}(t)=\sum_{\hat{p_{i}}\in\hat{\mathcal{P}}}w_{i}\min_{{p_{ti}}\in\mathcal{P}_{t}}{{\lVert p_{ti}-\hat{p_{i}}\rVert}^{2}_{2}},$
+Specifically, the weight $w_{i}$ for each $p_{i}$ is calculated as:
+$w_{i}=\frac{\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})}{\sum_{i}\exp(-{\lVert p_{i}-p_{r}\rVert}^{2}/2{r_{s}}^{2})},$ (2)
+where the support radius $r_{s}$ is set to $r/3$, and $r$ denotes to the patch radius. Then, the reconstruction loss at the $t$-th iteration can be defined as:
+${\mathcal{L}}_{Recon}(t)=\sum_{\hat{p_{i}}\in\hat{\mathcal{P}}}w_{i}\min_{{p_{ti}}\in\mathcal{P}_{t}}{{\lVert p_{ti}-\hat{p_{i}}\rVert}^{2}_{2}},$ (3)
 where $\hat{\mathcal{P}}$ denotes the predicted point cloud, $\mathcal{P}_{t}$ denotes the adaptive GT at the $t$-th iteration, and $p_{i}$ represents a point in point cloud $\mathcal{P}$.
-On the other hand, the rendering loss focuses on visually aligning the noisy point cloud with the adaptive GT using differentiable rendering techniques.
-As calculated by Eq.
-8, it primarily concentrates on optimizing the boundary regions of the point cloud:
-(8) ${\mathcal{L}}_{Render}(t)=\frac{1}{K}\sum_{i=1}^{K}\lvert V_{i}(\mathcal{P}_{t})-V_{i}(\hat{\mathcal{P}})\rvert,$
-where $\hat{\mathcal{P}}$ denotes the predicted point cloud, $\mathcal{P}_{t}$ denotes the adaptive GT at the $t$-th iteration, $K$ is the quantity of rendered views, and $V_{i}$ represents a rendered view in a specific camera position.
-Loss function of one iteration is a weighted combination of the reconstruction loss and the rendering loss, and the ultimate loss is obtained by aggregating loss contributions across iterations, which are summed up as Eq.
-9:
-(9) ${\mathcal{L}}=\sum_{t=1}^{T}({\mathcal{L}}_{Recon}(t)+\alpha{\mathcal{L}}_{Render}(t)),$
-where $T$ is number of iteration times, and weight $\alpha$ is 0.01 in experiments.
-5. Experiments
-5.1. Datasets and Implementation
-Table 1. Results on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
+Loss function of one iteration is a weighted combination of the reconstruction loss and the rendering loss, and the ultimate loss is obtained by aggregating loss contributions across iterations, which are summed up as:
+${\mathcal{L}}=\sum_{t=1}^{T}({\mathcal{L}}_{Recon}(t)+\alpha{\mathcal{L}}_{Render}(t)),$ (4)
+where $T$ is number of iteration times, and weight $\alpha$ is 0.01 in experiments. The detailed pseudocode for training 3DMambaIPF is provided in Appendix.
+Experiments
+Figure 5: Visualization Comparisons on Stanford 3D Scanning Repository with 5% Gaussian deviation.
+Coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in green, while those with high P2F values (noisy) are portrayed in purple.
+Figure 6: Visualization Comparisons on PU-Net Dataset of 50K points with 2.5% Gaussian deviation.
+The coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in blue, while those with high P2F values (noisy) are portrayed in red.
+Method Armadillo Dragon
+173K points 438K points
+CD P2M CD P2M
+Noisy 50.19 47.96 38.92 37.51
+Pointfilter 6.41 4.82 12.10 12.70
+PD-Flow 52.30 50.08 51.48 50.00
+ScoreDenoise 29.42 27.41 36.25 34.88
+IterativePFN 17.04 14.93 20.43 18.98
+3DMambaIPF (Ours) 17.14 15.06 19.79 18.36
+Table 1: Results on Stanford 3D Scanning Repository. CD and P2M distances are multiplied by $10^{5}$.
+Overview of Comparative Baselines
+The filtering capabilities of comparative baselines are conducted with DMRDenoise (Luo and Hu 2020), Pointfilter (Zhang et al.
+2020), PD-Flow (Mao et al.
+2022), ScoreDenoise (Luo and Hu 2021) and IterativePFN (de Silva Edirimuni et al.
+2023).
+IterativePFN and 3DMambaIPF work on highly-dense and large-scale datasets.
+Pointfilter tends to obtain over-smoothed surfaces that deviate from the GT and fail to keep the original structure after filtering and struggles to work on highly-dense point clouds.
+On the other hand, the point cloud filtering task is not addressed as a multi-modal issue in previous works.
+All the baselines focus solely on either the movement of individual noise points or the generation of clean points, which leads to cluttered visual results, particularly around the surface of the point clouds.
+Datasets and Experimental Settings
+Details about the PU-Net (Yu et al.
+2018) dataset can be found in Appendix.
+Due to the limited number of points in PU-Net dataset, which contains a maximum of 50K points, point clouds from the Stanford 3D Scanning Repository (Curless and Levoy 1996; Krishnamurthy and Levoy 1996) are utilized to evaluate the filtering performance of large-scale synthetic objects.
+The preprocessing of datasets is detailed in the Appendix.
+The experiments are conducted on NVIDIA A100 GPUs, and 3DMambaIPF is implemented with PyTorch 1.13.1 and CUDA 11.7.
+The Mamba modules are applied with 6 layers, and the total parameters of Mamba-Denoising Modules is 0.81M.
+Following previous works (Luo and Hu 2021; de Silva Edirimuni et al.
+2023), the quantitative indicators CD and P2M are adopted for evaluation and comparison.
+More details can be found in Appendix.
 Method 10K points 50K points
@@
 DMRDenoise 47.12 21.96 50.85 25.23 52.77 26.69 12.05 7.62 14.43 9.70 16.96 11.90
-PDFlow 21.26 6.74 32.46 13.24 36.27 17.02 6.51 4.16 12.70 9.21 18.74 14.26
+Pointfilter 24.61 7.30 35.34 11.55 40.99 15.05 7.58 4.32 9.07 5.07 10.99 6.29
+PD-Flow 21.26 6.74 32.46 13.24 36.27 17.02 6.51 4.16 12.70 9.21 18.74 14.26
 ScoreDenoise 25.22 7.54 36.83 13.80 42.32 19.04 7.16 4.00 12.89 8.33 14.45 9.58
-Pointfilter 24.61 7.30 35.34 11.55 40.99 15.05 7.58 4.32 9.07 5.07 10.99 6.29
 IterativePFN 20.55 5.01 30.43 8.43 33.53 10.46 6.05 3.02 8.03 4.36 10.15 5.88
 3DMambaIPF (Ours) 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
-Figure 6.
-Visualization Comparisons on PU-Net Dataset of 50K points with 2.5% Gaussian deviation.
-The coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in blue, while those with high P2F values (noisy) are portrayed in red.
+Table 2: Results on PU-Net Dataset. CD and P2M distances are multiplied by $10^{5}$.
+Rendering Views Mamba Layers 50K points
 […]
+CD P2M CD P2M CD P2M
+/ / 6.05 3.02 8.03 4.36 10.15 5.88
+/ 6 5.94 2.92 7.70 4.14 9.41 5.37
+8 6 6.04 2.99 8.06 4.36 10.39 6.05
+16 6 6.21 3.14 7.87 4.29 9.63 5.60
+24 6 5.96 2.97 7.81 4.26 9.56 5.53
+32 6 5.89 2.91 7.55 4.05 9.28 5.31
+32 1 8.15 4.71 10.82 6.53 13.23 8.20
+32 3 6.12 3.09 8.05 4.43 9.58 5.54
+32 6 5.89 2.91 7.55 4.05 9.28 5.31
+32 9 6.15 3.08 8.01 4.38 9.89 5.74
+32 12 6.12 3.07 8.09 4.42 9.85 5.67
+Table 3: Ablation results of Mamba layers and rendered views on PU-Net 50K dataset. CD and P2M distances are multiplied by $10^{5}$.
+Comparisons on Highly-dense and Large-scale Datasets
+The filtering performance of highly dense and large-scale point clouds is evaluated on synthetic models from the Stanford 3D Scanning Repository.
+The quantitative and visual results are shown in Table 1 and Figure 5.
+Due to the presence of dense and distant noise relative to the surface, some baselines struggle to effectively filter the point clouds.
+While Pointfilter achieves the best quantitative results, it lacks significant geometric details and accurate surface structures.
+It is found that the incorporation of the generative model into PD-Flow results in varied denoising outcomes across different experiments on the same point cloud.
+Despite potential variations in outcomes, PD-Flow generally falls short in effectively denoising highly-dense point clouds, particularly in severely noisy areas where it struggles to achieve significant noise reduction.
+ScoreDenoise, employing a patch partitioning method, exhibits gaps in results when processing highly-dense and large-scale point clouds.
+Although IterativePFN tackles this issue with patch-stitching and achieves competitive quantitative results, its performance is limited when dealing with point cloud surfaces and densely cluttered areas.
+As a single-modal filtering method, it struggles with the visual quality of point cloud surfaces, leading to gaps on the object surface that do not align with GT.
+On the contrary, 3DMambaIPF achieves the best performance when dealing with highly dense and large-scale point clouds.
+Additionally, the utilization of rendering loss results in significantly improved filtering effects and better preservation of geometric details, particularly around the surfaces of the point clouds.
+Comparisons on PU-Net Dataset
 […]
-Implementation.
-The 3DMambaIPF network is trained on NVIDIA A100 GPUs with PyTorch 1.13.1 and CUDA 11.7.
-The network underwent 100 epochs of training on PU-Net dataset, employing the Adam optimizer with a learning rate of 1 × $10^{-4}$.
-Comparative experiments are conducted with DMRDenoise (Luo and Hu, 2020), Pointfilter (Zhang et al., 2020), PD-Flow (Mao et al., 2022), ScoreDenoise (Luo and Hu, 2021) and IterativePFN (de Silva Edirimuni et al., 2023).
-All experimental methods (including 3DMambaIPF) are tested on the NVIDIA GeForce RTX 3090 GPU for consistency across evaluations in this paper.
-5.2. Evaluation metrics
-Figure 7.
-Visualization Comparisons on large-scale models from the Stanford 3D Scanning Repository with 5% Gaussian deviation.
-Coloration of each point is determined by its point-wise P2F distance, with points exhibiting low P2F values (clean) depicted in green, while those with high P2F values (noisy) are portrayed in purple.
-Figure 8. Visualization Comparisons on Paris-rue-Madame dataset.
-Table 2. Ablation results of loss function on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
-Method 10K points 50K points
-[…]
-CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
-IterativePFN 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
-3DMambaIPF (without view loss) 20.02 4.80 30.19 8.13 32.97 10.05 5.94 2.92 7.70 4.14 9.41 5.37
-3DMambaIPF (with view loss) 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
-Referring to many prior notable works, we utilize Chamfer Distance (CD) and Point-to-Mesh (P2M) Distance as the evaluation metrics for our study.
-Notably, CD is delineated in Eq.
-10, encompassing the combined sum of two components: the average distances between each point in the predicted point cloud and its nearest point in the GT point cloud, and the average distances between each point in the GT point cloud and its nearest point in the predicted point cloud.
-(10) ${\mathcal{L}}_{CD}(\hat{\mathcal{P}},\mathcal{P})=\frac{1}{|\hat{\mathcal{P}}|}\sum_{{\hat{p}\in\hat{\mathcal{P}}}}\min_{{p}\in\mathcal{P}}\|\hat{p}-{p}\|+\frac{1}{|\mathcal{P}|}\sum_{{p}\in\mathcal{P}}\min_{{\hat{p}\in\hat{\mathcal{P}}}}\|{p}-\hat{p}\|,$
-[…]
-(11) ${\mathcal{L}}_{P2M}(\hat{\mathcal{P}},\mathcal{M})={\mathcal{L}}_{P2F}(\hat{\mathcal{P}}$ $,\mathcal{M})+{\mathcal{L}}_{F2P}(\hat{\mathcal{P}},\mathcal{M}),$
-${\mathcal{L}}_{P2F}(\hat{\mathcal{P}},\mathcal{M})=\frac{1}{|\hat{\mathcal{P}}|}$ $\sum_{{\hat{p}\in\hat{\mathcal{P}}}}\min_{{f}\in\mathcal{M}}d(\hat{p},f),$
-${\mathcal{L}}_{F2P}(\hat{\mathcal{P}},\mathcal{M})=\frac{1}{|\mathcal{M}|}$ $\sum_{{f}\in\mathcal{M}}\min_{{\hat{p}\in\hat{\mathcal{P}}}}d(\hat{p},f),$
-where $\hat{\mathcal{P}}$ represents the predicted point cloud, $\mathcal{M}$ represents the mesh of GT point cloud, $|\mathcal{P}|$ signifies the number of points in $\mathcal{P}$, $|\mathcal{M}|$ signifies the number of faces in $\mathcal{M}$, and $d(p,f)$ denotes square distance from point $p$ to face $f$.
-Please note that due to the typically unequal quantities of faces and points, point-wise calculation of P2M distance is unreasonable.
-In the visualization of denoising results in Section 5.3 and Section 5.4, we utilize the point-wise P2F distance from Eq.
-11 to visualize the denoised point cloud results.
-Table 3. Ablation results of views in point rendering on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
-Method Rendered Views 10K points 50K points
-[…]
-CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
-IterativePFN / 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
-3DMambaIPF 8 20.36 4.96 30.73 8.48 34.35 10.96 6.04 2.99 8.06 4.36 10.39 6.05
-3DMambaIPF 16 20.29 4.97 30.35 8.24 33.08 10.21 6.21 3.14 7.87 4.29 9.63 5.60
-3DMambaIPF 24 19.96 4.82 30.26 8.14 32.85 9.95 5.96 2.97 7.81 4.26 9.56 5.53
-3DMambaIPF 32 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
-Table 4. Ablation results of Mamba layers on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
-Method Mamba Layers 10K points 50K points
-[…]
-CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
-IterativePFN / 20.56 5.01 30.43 8.45 33.52 10.45 6.05 3.02 8.03 4.36 10.15 5.88
-3DMambaIPF 1 23.35 6.76 33.24 10.91 36.42 13.17 8.15 4.71 10.82 6.53 13.23 8.20
-3DMambaIPF 3 20.24 5.02 30.59 8.61 33.44 10.60 6.12 3.09 8.05 4.43 9.58 5.54
-3DMambaIPF 6 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
-3DMambaIPF 9 20.55 5.11 30.77 8.59 33.82 10.70 6.15 3.08 8.01 4.38 9.89 5.74
-3DMambaIPF 12 20.48 5.05 30.83 8.80 33.92 10.98 6.12 3.07 8.09 4.42 9.85 5.67
-Table 5. Ablation results of patch size on PU-Net Dataset. CD and P2M distances are multiplied by $\textbf{10}^{\textbf{5}}$.
-Method Patch Size 10K points 50K points
-[…]
-CD P2M CD P2M CD P2M CD P2M CD P2M CD P2M
-IterativePFN 1000 20.55 5.01 30.43 8.43 33.53 10.46 6.05 3.02 8.03 4.36 10.15 5.88
-3DMambaIPF (Ours) 1000 19.89 4.77 29.95 8.03 32.62 9.92 5.89 2.91 7.55 4.05 9.28 5.31
-IterativePFN 2000 20.64 5.02 30.43 8.41 33.93 10.72 6.01 2.99 7.79 4.19 10.25 5.95
-3DMambaIPF (Ours) 2000 19.98 4.81 30.05 8.14 32.73 10.04 5.92 2.92 7.52 4.05 8.98 5.13
-5.3. Comparisons on Small-Scale Datasets
-[…]
-5.4. Comparisons on Large-Scale Datasets
-The visualization performance of large-scale point clouds is tested on both large-scale synthetic models from the Stanford 3D Scanning Repository and Paris-rue-Madame dataset.
-In the visualization results of large-scale synthetic datasets, some of the baselines struggle to filter due to the presence of dense and distant noise from the surface.
-While PointFilter achieves filtering results where each point is close to a clean surface, it lacks significant geometric details and accurate edge structures.
-It is found that the incorporation of the generative model into PD-Flow results in varied denoising outcomes across different experiments on the same point cloud.
-Despite potential variations in outcomes, PD-Flow generally falls short in effectively denoising large-scale point clouds, particularly in densely noisy areas where it struggles to achieve significant noise reduction.
-ScoreDenoise, employing a patch partitioning method, exhibits gaps in results when processing large-scale point clouds.
-Although IterativePFN addresses this issue with patch-stitching, its performance suffers when handling point cloud edges and densely cluttered point clouds.
-On the contrary, the inclusion of the Mamba module enhances the processing capabilities of 3DMambaIPF when dealing with dense point clouds.
-Additionally, the utilization of rendering loss results in remarkable denoising effects specifically on the edges of the point cloud.
-In the visualization of real-world large-scale dataset (Paris-rue-Madame), when there are no additional noises introduced, all the baseline methods are capable of achieving satisfactory filtering results.
-3DMambaIPF outperforms the others in terms of preserving geometric edges.
-6. Ablation Study
-Ablation experiments are conducted on the influencing factors of 3DMambaIPF, focusing on four aspects: i) loss function, ii) the quantity of rendered images in point rendering, iii) the number of Mamba layers, and iv) patch size.
-All the ablation experiments are tested on PU-Net dataset with NVIDIA GeForce RTX 3090 GPU.
-Initially, Table 2 validates the enhancement of the filtering capability of 3DMambaIPF across various levels of noise with the introduction of view loss.
-Following the previous validation step, we proceed to examine how varying numbers of rendered views affect filtering effectiveness.
-Table 3 suggests that generally, a greater number of views correlates with stronger denoising ability, with 32 views yielding the best results.
-However, employing an excessive number of views will decrease the training speed.
+Ablation Study
+Ablation on Modules
+Ablation experiments are firstly conducted on the influencing factors of 3DMambaIPF, focusing on two aspects: i) the quantity of rendered images in point rendering, ii) the number of Mamba layers.
+Table 3 validates the enhancement of the filtering capability of 3DMambaIPF across various levels of noise with the introduction of view loss, and proceeds to examine how varying numbers of rendered views affect filtering effectiveness.
+However, employing an excessive number of views will decrease the training speed, and detailed runtime metrics will be validated in appendix.
 Subsequently, ablation experiments are carried out on the number of Mamba layers.
 Table 3 illustrates that the denoising capability is at its peak with 6 Mamba layers.
-At last, the influence of patch size on denoising capability is examined in Table 5, finding that larger patch sizes are more suitable for larger scenes.
-7. Conclusion
-In this study, we introduce 3DMambaIPF, an iterative point cloud filtering model that utilizes the Mamba module.
-3DMambaIPF not only achieves state-of-the-art results on small-scale (PU-Net) dataset, but also demonstrates superior performance on large-scale datasets, indicating its capability to address denoising challenges in extensive scenes, as validated through ablation experiments.
+Runtime of Differentiable Rendering
+An additional ablation for the runtime of rendered views is conducted as shown in Table 4.
+RuntimeDR represents the total time required to render both the GT and the predicted point clouds.
+RuntimeStep denotes the total time for one training step.
+Due to our devised network-free differentiable rendering, the runtime of differentiable rendering only utilizes 11.8%-16.5% of the total time for one training step.
+Except for this efficiency, differentiable rendering demonstrates notable improvements of 12.0% and 13.9% in CD and P2M metrics, respectively, as the number of views increases.
+This trade-off between runtime and performance is cost-effective.
+Views Metrics
+CD P2M RuntimeDR (s) RuntimeStep (s)
+8 10.39 6.05 0.08 0.74
+16 9.63 5.60 0.11 0.82
+24 9.56 5.53 0.14 0.97
+32 9.28 5.31 0.17 1.03
+Table 4: Runtime analysis on differentiable rendering. CD and P2M metrics are multiplied by $10^{5}$, and tested on PU-Net Dataset of 50K points with 2.5% Gaussian deviation.
+Comparisons 3D Mamba with Other Backbones
+To validate the effectiveness of our designed 3D Mamba, ablation studies (Table 5) comparing the Mamba modules with Transformers and MLPs are conducted.
+For fairness, the number of modules, layers and embedding sizes are consistent with 3DMambaIPF.
+For intuitive results, differentiable rendering loss is not implemented in this ablation.
+3DMambaIPF fulfills the best denoising performance while ensuring fast training speed as well as the smallest model size.
+Methods Metrics
+CD P2M GFLOPs FPS MParams
+with Transformers 14.02 8.56 178.76 2.31 158.00
+with MLPs 30.17 21.71 20.78 4.39 4.56
+with Mambas (Ours) 9.42 5.37 16.92 3.62 0.81
+Table 5: Comparsions on Mamba with other backbones. CD and P2M metrics are multiplied by $10^{5}$, and tested on PU-Net Dataset of 50K points with 2.5% Gaussian deviation.
+Conclusion
+In this study, we introduce 3DMambaIPF, a novel iterative Mamba-based point cloud filtering model that utilizes a differentiable rendering technique for the first time in this area.
+3DMambaIPF not only achieves state-of-the-art results on a commonly-used dataset but also demonstrates superior performance on a highly dense and large-scale dataset, indicating its capability to address denoising challenges in extensive scenes, as validated through ablation experiments.
 Although 3DMambaIPF may encounter reverse-amplification issues on denoising a smooth and clean surface, its performance remains optimal.
+In future work, we will explore different datasets to address the challenges of filtering real-world point clouds from large scenes.
+Acknowledgments
+This project was partially supported by the National Natural Science Foundation of China (U2033209); Hong Kong Innovation and Technology Fund (PiH/213/24GS); the Ministry of Education, Singapore, under its Academic Research Fund Grants (MOE-T2EP20220-0005 & RT19/22) and an OPPO gift fund.